# Approximating the set of Nash equilibria for convex games

### v2

@@ 4.4 Algorithm and main result
Compute an inner approximation $\mathcal{P}^{In}_{i}$ such that (10) is satisfied with approximation level $\varepsilon_{1}>0$ using [11, Algorithm 1] with the modified initialization using (11) and (12) as described in Section 4.1.
@@
This can be done by e.g. [19, 12, 8] or [9, Algorithm 5.1], see also Section 4.3. Then, $\hat{X}^{j}_{i}:=\{x\in\mathbb{X}\;|\;(x,y)\in\bar{S}\}$ yields an $\varepsilon_{2}$-approximation of the set $\bar{X}^{j}_{i}$ in the sense of (15).
Set
@@
We can now state the main theorem of this paper.
@@
Let Assumptions 4.1 and 4.2 be satisfied. Let $X\subseteq\mathbb{R}^{\sum_{i=1}^{N}n_{i}}$ be the set computed by Algorithm 4.7. Then it holds
$\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$ (16)
for $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$.
@@
Let $\mathcal{P}^{In}_{i}=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$ for $\bar{\mathcal{X}}=\{{x}^{1},...,{x}^{s}\}\subseteq\mathbb{X}$ with $s\in\mathbb{N}$ be the inner approximation of $\mathcal{P}_{i}$ computed as described in Section 4.1. Due to Remark 4.5, we know there are $\lambda^{1},...,\lambda^{s}\geq 0$ with $\sum_{l=1}^{s}\lambda^{l}=1$ such that $x^{*}_{-i}=\sum_{l=1}^{s}\lambda^{l}{x}^{l}_{-i}$. Thus, $(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))\in\mathcal{P}^{In}_{i}$. Then there is $\hat{y}$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})$ is minimal in $\mathcal{P}^{In}_{i}$: Assume there would be no such $\hat{y}$. Then for all $\alpha>0$ it would be $(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))-\alpha\bar{c}_{i}\in\mathcal{P}^{In}_{i}$. In that case $\mathcal{P}^{In}_{i}$ would contain the line $\{(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))+\alpha\bar{c}_{i}\;|\;\alpha\in\mathbb{R}\}$ which is a contradiction to $\mathcal{P}^{In}_{i}$ being pointed. Thus, there exists $\hat{y}$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})$ is minimal in $\mathcal{P}^{In}_{i}$ and therefore exists a maximal efficient face $F^{j}$ of $\mathcal{P}^{In}_{i}$ with $(x^{*}_{-i},-x^{*}_{-i},\hat{y})\in F^{j}$. So there are $\hat{\lambda}^{1},...,\hat{\lambda}^{s}\geq 0$ with $\sum_{l=1}^{s}\hat{\lambda}^{l}=1$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})=(\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}_{-i},-\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}_{-i},\sum_{l=1}^{s}\hat{\lambda}^{l}f_{i}({x}^{l}))$. Since $f_{i}(x^{*})\leq f_{i}(\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l})\leq\hat{y}$ it is $x^{*}\in\bar{X}^{j}_{i}$.
@@
Let Assumptions 4.1 and 4.2 be satisfied. For any $i=1,...,N$ it holds
$\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}\subseteq X_{i}\subseteq\epsilon\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$
for $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$.
@@
Let us first prove the left hand side inclusion. Let $x^{*}\in\operatorname{NE}(f,\mathbb{X})$. By Theorem 3.1, it follows that $x^{*}\in\bigcap_{i=1}^{N}\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$. Then, for each $i=1,...,N$ there is by Lemma 4.9 a $j=j(i)$ such that $x^{*}\in\bar{X}^{j}_{i}$. This implies, by (15), $x^{*}\in\operatorname{conv}\hat{X}^{j}_{i}+B_{\varepsilon_{2}}$. As $x^{*}\in\operatorname{NE}(f,\mathbb{X})$ also implies $x^{*}\in\mathbb{X}$, this proves the claim by the definition of the sets $X_{i}$ and $X$. Let us now prove the right hand side inclusion. Let $x^{*}\in X$, thus $x^{*}\in X_{i}$ for all $i=1,...,N$. By Lemma 4.10 and Theorem 3.2 the claim follows.
@@
Theorem 4.8 also holds if we replace Assumption 4.2 by the following condition, keeping Assumption 4.1 unchanged: Assume each player $i$ has an independent constraint set $\mathbb{X}_{i}=\{x_{i}\in\mathcal{X}_{i}\;|\;g_{i}(x_{i})\leq 0\}$, where $g_{i}:\mathcal{X}_{i}\to\mathbb{R}$ is a convex function and $\mathbb{X}_{i}$ has nonempty interior. Thus, the shared constraint set is $\mathbb{X}=\mathbb{X}_{1}\times...\times\mathbb{X}_{N}=\{x\in\prod_{j=1}^{N}\mathcal{X}_{j}\;|\;g_{1}(x_{1})\leq 0,...,g_{N}(x_{N})\leq 0\}$. Constraints of this type are also frequently considered in the literature: they are called ‘orthogonal constraint sets’ in [18] and ‘classical games’ in [2].
When replacing the polyhedral constraint Assumption 4.2 by the independent constraint assumption $\mathbb{X}=\mathbb{X}_{1}\times...\times\mathbb{X}_{N}$, one needs to adapt step $1$ of Algorithm 4.7 as follows
@@
Compute an inner approximation $\mathcal{P}^{In}_{i}$ of its upper image such that (10) is satisfied with approximation level $\varepsilon_{1}>0$ using [11, Algorithm 1] with the modified initialization, similar to Section 4.1, but using
@@
The rest of Algorithm 4.7 remains unchanged. The set $X$ outputted by this modification of Algorithm 4.7 will satisfy
$\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X}),$
for $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$ for a game with an independent constraint set as above and satisfying Assumption 4.1. The proofs are in analogy to the proof of Theorem 4.8. The key is that the independent constraint assumption and the above stated modification of step $1$ of Algorithm 4.7 ensure that $\mathcal{P}_{i}^{In}$ has the same property as in Remark 4.5.
@@
Note however, that for a more general constraint set $\mathbb{X}$ than that considered in Assumption 4.2 or Remark 4.11, the property stated in Remark 4.5 is not enough to prove that Algorithm 4.7 outputs a set $X$ that satisfies (16). In the independent constraint case of Remark 4.11, the set $\bar{\mathcal{X}}:=\{x^{1},...,x^{s}\}$ that is computed in step 1 of the algorithm has for each $i$ the following two properties:
  • $\bar{\mathcal{X}}_{-i}:=\operatorname{conv}\{x^{1}_{-i},...,x^{s}_{-i}\}\supseteq\mathbb{X}_{-i}:=\{x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}\;|\;\exists\;x_{i}\in\mathcal{X}_{i}:(x_{i},x_{-i})\in\mathbb{X}\}$
  • $\bar{\mathcal{X}}_{i}:=\operatorname{conv}\{x^{1}_{i},...,x^{s}_{i}\}\subseteq\mathbb{X}_{i}:=\{x_{i}\in\mathcal{X}_{i}\;|\;\exists\;x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}:(x_{i},x_{-i})\in\mathbb{X}_{i}\}$.
Due to the independent constraint sets, one obtains that for each $\tilde{x}_{i}\in\bar{\mathcal{X}}_{i}$ and for each $\hat{x}_{-i}\in\bar{\mathcal{X}}_{-i}$ it holds $(\tilde{x}_{i},\hat{x}_{-i})\in\mathbb{X}$, which is needed in the proof of Lemma 4.9 under the assumption of Remark 4.11. However, this property cannot be guaranteed when we would add constraints that have no independence for the different strategies.
@@ 5 Examples
We will start in Section 5.1 with simple illustrative two player examples where the set of Nash equilibria, as well as the set of $\epsilon$-approximate Nash equilibria, can be computed explicitly. As such we directly investigate the sandwich principle (16). In Section 5.2, we will consider more complex examples from environmental economics involving two or three players. There, the set of Nash equilibria or $\epsilon$-approximate Nash equilibria is in general not known and Algorithm 4.7 is used to compute a set $X$ that by Theorem 4.8 is known to lie between the set of true Nash equilibria and the set of $\epsilon$-approximate Nash equilibria.
@@ 5.1 Illustrative examples
Consider a two player game, where each player chooses a real-valued strategy, thus $\mathcal{X}_{1}=\mathcal{X}_{2}=\mathbb{R}$. Let the shared constraint set be $\mathbb{X}=\{(x_{1},x_{2})\in\mathbb{R}^{2}\;|\;x_{1},x_{2}\in[-1,1]\}$. Clearly, the set $\mathbb{X}$ is a compact polyhedral set. Consider for player $1$ the convex cost function $f_{1}(x_{1},x_{2})=\frac{1}{2}x^{2}_{1}-x_{1}(x_{2}+\frac{1}{2})+x^{2}_{2}$ and for player $2$ the convex cost $f_{2}(x_{1},x_{2})=\frac{1}{2}x^{2}_{2}+x_{1}x_{2}+x^{2}_{1}$. One can show that $f_{1}$ and $f_{2}$ are Lipschitz continuous with constant $L=3$ on $\mathbb{X}$. Thus, Assumptions 4.1 and 4.2 are satisfied. Due to the simple structure and low dimension of this example it is possible to find the set of Nash equilibria easily by hand. It consists of the unique Nash equilibrium $(\frac{1}{4},-\frac{1}{4})$. However, we will apply Algorithm 4.7 to illustrate the sandwich result (16) of Theorem 4.8 and thus find an approximation of the unique Nash equilibrium. We choose error levels $\varepsilon_{1}=0.01,\varepsilon_{2}=0.001$. Hence, the overall error is $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}=0.016$. Due to the simple nature of the problem we are also able to formulate the exact set of $\epsilon$-approximate Nash equilibria for this example, which is given by
@@
Figure 1: Computed sets $X_{1}$ for player $1$ (blue) and $X_{2}$ for player $2$ (yellow) for Example 5.1. The intersection for both players yields the set $X$ that satisfies $\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$ for $\epsilon=0.016$.
Figure 2: The singleton set $\operatorname{NE}(f,\mathbb{X})$ (red), the set $X$ computed by Algorithm 4.7 (blue) and the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ (yellow) for Example 5.1.
@@
Consider now the example provided within [4, Section 2]. This is a parametric game which permits unique, multiple, and infinite Nash equilibria depending on the parameter chosen.
@@
Figure 3: Computed sets $X_{1}$ for player $1$ (blue) and $X_{2}$ player $2$ (yellow) for (a) $y=0.5$, (b) $y=1$ and (c) $y=1.5$ in Example 5.2.
@@
In order to apply Algorithm 4.7, we consider the compact subset $\bar{\mathbb{X}}=\mathbb{X}\cap[0,2]^{2}$. Since $\operatorname{NE}(f,\mathbb{X})\subseteq\bar{\mathbb{X}}$, there is no loss of information by using this smaller constraint set for the computations. The Lipschitz constant of the joint cost function $(f_{1},f_{2})$ over $\bar{\mathbb{X}}$ is $L=8$. We apply Algorithm 4.7 for chosen error levels $\varepsilon_{1}=0.01,\varepsilon_{2}=0.001$. Thus, the outputted set $X\subseteq\mathbb{R}^{2}$ satisfies (16) for $\epsilon=0.027$. Figure 5 shows the computed sets $X_{1}$ and $X_{2}$ satisfying (9) for both players and its intersection $X$.
Figure 5: Example 5.3: Computed sets $X_{1}$ for player 1 (blue) and $X_{2}$ for player 2 (yellow). The intersection for both players yields the set $X$ that satisfies $\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$ for $\epsilon=0.027$.
The exact set of $\epsilon$-approximate Nash equilibria for this example is given by
@@
for $v_{1}(x_{2})=\frac{1}{2}(x_{2}\vee(1-x_{2}))^{2}-(x_{2}\vee(1-x_{2}))x_{2}$ and $v_{2}(x_{1})=(1-x_{1})^{+}$. Figure 6 shows the three sets of the sandwich result $\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$. Similar to before, the computed set $X$ is much smaller than the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ and approximates the set of true Nash equilibria $\operatorname{NE}(f,\mathbb{X})$ quite well.
Figure 6: Example 5.3: the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ (yellow), the computed set $X$ (blue) and the set $\operatorname{NE}(f,\mathbb{X})$ (red) as well as a zoomed in region.

### v4

@@ 4.4 Algorithm and main result
Compute an inner approximation $\mathcal{P}^{In}_{i}$ such that (10) is satisfied with approximation level $\varepsilon_{1}>0$ using [14, Algorithm 1] with the modified initialization using (11) and (12) as described in Section 4.1 or using [14, Algorithm 2] with modification (12) as discussed in Remark 4.6.
@@
This can be done by e.g. [22, 15, 10] or [11, Algorithm 5.1], see also Section 4.3. Then, $\hat{X}^{j}_{i}:=\{x\in\mathbb{X}\;|\;(x,y)\in\bar{S}\}$ yields an $\varepsilon_{2}$-approximation of the set $\bar{X}^{j}_{i}$ in the sense of (15). Set
@@
We can now state the main theorem of this paper. Recall that $L>0$ is the largest of the Lipschitz constants of the functions $f_{i}$.
@@
Let Assumptions 4.1 and 4.2 be satisfied. Let $X\subseteq\mathbb{R}^{\sum_{i=1}^{N}n_{i}}$ be the set computed by Algorithm 4.8 and let $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$. Then it holds
$\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X}).$ (16)
@@
Let $\mathcal{P}^{In}_{i}=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$ for $\bar{\mathcal{X}}=\{{x}^{1},...,{x}^{s}\}\subseteq\mathbb{X}$ with $s\in\mathbb{N}$ be the inner approximation of $\mathcal{P}_{i}$ computed as described in Section 4.1. Due to Remark 4.5, we know there are $\lambda^{1},...,\lambda^{s}\geq 0$ with $\sum_{l=1}^{s}\lambda^{l}=1$ such that $x^{*}_{-i}=\sum_{l=1}^{s}\lambda^{l}{x}^{l}_{-i}$. Thus, $(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))\in\mathcal{P}^{In}_{i}$. Then there is $\hat{y}$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})$ is minimal in $\mathcal{P}^{In}_{i}$: Assume there would be no such $\hat{y}$. Then for all $\alpha>0$ it would be $(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))-\alpha\bar{c}_{i}\in\mathcal{P}^{In}_{i}$. In that case $\mathcal{P}^{In}_{i}$ would contain the line $\{(x^{*}_{-i},-x^{*}_{-i},\sum_{l=1}^{s}\lambda^{l}f_{i}({x}^{l}))+\alpha\bar{c}_{i}\;|\;\alpha\in\mathbb{R}\}$ which is a contradiction to $\mathcal{P}^{In}_{i}$ being pointed. Thus, there exists $\hat{y}$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})$ is minimal in $\mathcal{P}^{In}_{i}$ and therefore exists a maximal efficient face $F^{j}$ of $\mathcal{P}^{In}_{i}$ with $(x^{*}_{-i},-x^{*}_{-i},\hat{y})\in F^{j}$. So there are $\hat{\lambda}^{1},...,\hat{\lambda}^{s}\geq 0$ with $\sum_{l=1}^{s}\hat{\lambda}^{l}=1$ such that $(x^{*}_{-i},-x^{*}_{-i},\hat{y})=(\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}_{-i},-\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}_{-i},\sum_{l=1}^{s}\hat{\lambda}^{l}f_{i}({x}^{l}))$. Since
${}f_{i}(x^{*})\leq f_{i}(\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l})\leq\sum_{l=1}^{s}\hat{\lambda}^{l}f_{i}({x}^{l})=\hat{y},$ (17)
by $x^{*}\in\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$, $\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}\in\mathbb{X}$, and $f_{i}$ convex, it is $x^{*}\in\bar{X}^{j}_{i}$.
@@
Let Assumptions 4.1 and 4.2 be satisfied. Let $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$. For any $i=1,...,N$ it holds
$\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}\subseteq X_{i}\subseteq\epsilon\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}.$
@@
Let us first prove the left hand side inclusion. Let $x^{*}\in\operatorname{NE}(f,\mathbb{X})$. By Theorem 3.1, it follows that $x^{*}\in\bigcap_{i=1}^{N}\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$. Then, for each $i=1,...,N$ there is by Lemma 4.10 a $j=j(i)$ such that $x^{*}\in\bar{X}^{j}_{i}$. This implies, by (15), $x^{*}\in\operatorname{conv}\hat{X}^{j}_{i}+B_{\varepsilon_{2}}$. As $x^{*}\in\operatorname{NE}(f,\mathbb{X})$ also implies $x^{*}\in\mathbb{X}$, this proves the claim by the definition of the sets $X_{i}$ and $X$. Let us now prove the right hand side inclusion. Let $x^{*}\in X$, thus $x^{*}\in X_{i}$ for all $i=1,...,N$. By Lemma 4.11 and Theorem 3.2 the claim follows. ∎
@@
Recall that Theorem 3.1 was given in [6] in a more general framework and in a different formulation using a non-standard ordering cone (see problem (5) with ordering cone (7) in [6, Cor. 2.8]), which can be equivalently written as
$\mathbb{R}_{+}\bar{c}_{i}-\min\{(x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\},$ (18)
where one is minimizing just in direction $\bar{c}_{i}=(0,...,0,1)^{\top}\in\mathbb{R}^{\sum_{j=1,j\neq i}^{N}n_{j}+1}$. That is, the ordering cone of problem (18) is not the natural ordering cone, but $\mathbb{R}_{+}\bar{c}_{i}=\{r\bar{c}_{i}\;|\;r\in\mathbb{R}_{+}\}$.
Thus, instead of problem (3) one could also consider problem (18), which is lower dimensional (one dimension lower than the dimension reduction obtained already by (4)), but has a non-standard ordering cone. One can, however, not expect a significant run time reduction of the proposed algorithm for this lower dimensional problem as the scalar problems solved within the algorithms in step 1 and 3 of the Algorithm 4.8 are basically unaffected by this change. Step 2 would be in lower dimensions, but the run time bottle neck of Algorithm 4.8 are rather steps 1 and 3.
Remark 4.13.
Theorem 4.9 also holds if we replace Assumption 4.2 by the following condition, keeping Assumption 4.1 unchanged: Assume each player $i$ has an independent constraint set $\mathbb{X}_{i}=\{x_{i}\in\mathcal{X}_{i}\;|\;g_{i}(x_{i})\leq 0\}$, where $g_{i}:\mathcal{X}_{i}\to\mathbb{R}$ is a convex function and $\mathbb{X}_{i}$ has nonempty interior. Thus, the shared constraint set is $\mathbb{X}=\mathbb{X}_{1}\times...\times\mathbb{X}_{N}=\{x\in\prod_{j=1}^{N}\mathcal{X}_{j}\;|\;g_{1}(x_{1})\leq 0,...,g_{N}(x_{N})\leq 0\}$. Constraints of this type are also frequently considered in the literature: they are called ‘orthogonal constraint sets’ in [21] and ‘classical games’ in [2]. When replacing the polyhedral constraint Assumption 4.2 by the independent constraint assumption $\mathbb{X}=\mathbb{X}_{1}\times...\times\mathbb{X}_{N}$, one needs to adapt step $1$ of Algorithm 4.8 as follows
@@
Compute an inner approximation $\mathcal{P}^{In}_{i}$ of its upper image such that (10) is satisfied with approximation level $\varepsilon_{1}>0$ using [14, Algorithm 1] with the modified initialization, similar to Section 4.1, but using
@@
The rest of Algorithm 4.8 remains unchanged. The set $X$ outputted by this modified Algorithm 4.8 satisfies
$\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$
for $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}$ for a game with an independent constraint set as above and satisfying Assumption 4.1. The proofs are in analogy to the proof of Theorem 4.9. The key is that the independent constraint assumption ($\mathbb{X}_{i}\times\mathbb{X}_{-i}=\mathbb{X}$) ensures that the first inequality in (17) still holds by $x^{*}\in\operatorname*{argMin}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$ and since $\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}=(\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}_{i},x^{*}_{-i})\in\mathbb{X}_{i}\times\mathbb{X}_{-i}=\mathbb{X}$ is feasible, as ${x}^{l}_{i}\in\mathbb{X}_{i}$ ($l=1,...,s$) even though ${x}^{l}\notin\mathbb{X}$ is possible due to the use of the outer approximation $P^{Out}_{-i}$ for $\mathbb{X}_{-i}$.
The approximation error of the outer approximation $P^{Out}_{-i}$ of $\mathbb{X}_{-i}$ does not enter as, in step $3$ of the algorithm, the intersection with the constraint set $\mathbb{X}$ is taken.
@@
Note however, that for a more general constraint set $\mathbb{X}$ than that considered in Assumption 4.2 or Remark 4.13, one might have that $\mathbb{X}$ is a strict subset of $\mathbb{X}_{i}\times\mathbb{X}_{-i}$, i.e. $\mathbb{X}\subsetneq\mathbb{X}_{i}\times\mathbb{X}_{-i}$, where we define $\mathbb{X}_{-i}:=\{x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}\;|\;\exists\;x_{i}\in\mathcal{X}_{i}:(x_{i},x_{-i})\in\mathbb{X}\}$ and $\mathbb{X}_{i}:=\{x_{i}\in\mathcal{X}_{i}\;|\;\exists\;x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}:(x_{i},x_{-i})\in\mathbb{X}\}$. Therefore the first inequality in (17) might fail as $\sum_{l=1}^{s}\hat{\lambda}^{l}{x}^{l}\notin\mathbb{X}$ is possible.
@@ 5 Examples
We will start in Sections 5.1 and 5.2 with simple illustrative two player examples where the set of Nash equilibria, as well as the set of $\epsilon$-approximate Nash equilibria, can be computed explicitly. As such we directly investigate the sandwich principle (16). In Section 5.3, we will consider more complex examples from environmental economics involving two or three players. There, the set of Nash equilibria or $\epsilon$-approximate Nash equilibria is in general not known and Algorithm 4.8 is used to compute a set $X$ that by Theorem 4.9 is known to lie between the set of true Nash equilibria and the set of $\epsilon$-approximate Nash equilibria.
@@ 5.1 Illustrative examples
Consider a two player game, where each player chooses a real-valued strategy, thus $\mathcal{X}_{1}=\mathcal{X}_{2}=\mathbb{R}$. Let the shared constraint set be $\mathbb{X}=\{(x_{1},x_{2})\in\mathbb{R}^{2}\;|\;x_{1},x_{2}\in[-1,1]\}$. Clearly, the set $\mathbb{X}$ is a compact polyhedral set. Consider for player $1$ the convex cost function $f_{1}(x_{1},x_{2})=\frac{1}{2}x^{2}_{1}-x_{1}(x_{2}+\frac{1}{2})+x^{2}_{2}$ and for player $2$ the convex cost $f_{2}(x_{1},x_{2})=\frac{1}{2}x^{2}_{2}+x_{1}x_{2}+x^{2}_{1}$. One can show that $f_{1}$ and $f_{2}$ are Lipschitz continuous with constant $L=3$ on $\mathbb{X}$. Thus, Assumptions 4.1 and 4.2 are satisfied. Due to the simple structure and low dimension of this example it is possible to find the set of Nash equilibria easily by hand. It consists of the unique Nash equilibrium $(\frac{1}{4},-\frac{1}{4})$. However, we will apply Algorithm 4.8 to illustrate the sandwich result (16) of Theorem 4.9 and thus find an approximation of the unique Nash equilibrium. We choose error levels $\varepsilon_{1}=0.01,\varepsilon_{2}=0.001$. Hence, the overall error is $\epsilon=\varepsilon_{1}+2L\varepsilon_{2}=0.016$. Due to the simple nature of the problem one can compute the exact set of $\epsilon$-approximate Nash equilibria as
@@
Figure 1: (a): Computed sets $X_{1}$ for player $1$ (blue) and $X_{2}$ for player $2$ (yellow), (b): The singleton set $\operatorname{NE}(f,\mathbb{X})$ (red), the set $X$ computed by Algorithm 4.8 (blue) and the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ (yellow) for Example 5.1.
@@
Consider now the example provided within [5, Section 2]. This is a parametric game which permits unique, multiple, and infinite Nash equilibria depending on the parameter chosen.
@@
Figure 2: Computed sets $X_{1}$ for player $1$ (blue) and $X_{2}$ for player $2$ (yellow) for (a) $y=0.5$, (b) $y=1$ and (c) $y=1.5$ in Example 5.2.
@@
In order to apply Algorithm 4.8, we consider the compact subset $\bar{\mathbb{X}}=\mathbb{X}\cap[0,2]^{2}$. Since $\operatorname{NE}(f,\mathbb{X})\subseteq\bar{\mathbb{X}}$, there is no loss of information by using this smaller constraint set for the computations. The Lipschitz constant of the joint cost function $(f_{1},f_{2})$ over $\bar{\mathbb{X}}$ is $L=8$. We apply Algorithm 4.8 for chosen error levels $\varepsilon_{1}=0.01,\varepsilon_{2}=0.001$. Thus, the outputted set $X\subseteq\mathbb{R}^{2}$ satisfies (16) for $\epsilon=0.027$. Figure 4 (a) shows the computed sets $X_{1}$ and $X_{2}$ satisfying (9) for both players and its intersection $X$. The exact set of $\epsilon$-approximate Nash equilibria for this example is given by
@@
for $v_{1}(x_{2})=\frac{1}{2}(x_{2}\vee(1-x_{2}))^{2}-(x_{2}\vee(1-x_{2}))x_{2}$ and $v_{2}(x_{1})=(1-x_{1})^{+}$. Figure 4 (b) shows the three sets of the sandwich result $\operatorname{NE}(f,\mathbb{X})\subseteq X\subseteq\epsilon\operatorname{NE}(f,\mathbb{X})$. Similar to before, the computed set $X$ is much smaller than the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ and approximates the set of true Nash equilibria $\operatorname{NE}(f,\mathbb{X})$ quite well.
Figure 4: Example 5.3: (a) Computed sets $X_{1}$ for player 1 (blue) and $X_{2}$ for player 2 (yellow). (b): the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ (yellow), the computed set $X$ (blue) and the set $\operatorname{NE}(f,\mathbb{X})$ (red) as well as a zoomed in region.
5.2 Search model
Consider the $N$-player search model from [3] and as described in [26, Section 3.1]. In this game, player $i$ spends effort $x_{i}$ (chosen from the space $\mathcal{X}_{i}=\mathbb{R}$) to find a partner for a transaction; each player is constrained so as to only allow for nonnegative efforts $\mathbb{X}\subseteq\mathbb{R}^{N}_{+}$. The cost of this effort is provided by a smooth and strictly convex function $C:\mathbb{R}_{+}\to\mathbb{R}_{+}$ that depends only on the action of the player, but is homogenous for all players. The expected payout is proportional to the effort put in by the player and some function $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ of the aggregate effort of all other players. Each player seeks to minimize her net costs $f_{i}(x):=C(x_{i})-x_{i}g(\sum_{j\neq i}x_{j})$.
Remark 5.4.
In [26, Section 3.1], no constraints are imposed on this game. Herein, so as to satisfy Assumption 4.1, we impose the constraint set $\mathbb{X}=[0,1]^{N}$. In addition, to guarantee joint convexity of the cost function, we consider the modified cost functions $\hat{f}_{i}(x)=f_{i}(x)+\beta\sum_{j\neq i}x_{j}^{2}$ with $\beta>0$ large enough.
Example 5.5.
Consider the case of two players. Assume a cubic cost function $C(z):=z^{3}$ and linear payout $g(z):=0.5z$. It is trivial to verify that there exist two Nash equilibria for this problem $\{(0,0)\,,\,(1/6,1/6)\}$. To satisfy the conditions of Assumption 4.1 we include the two modifications to this original problem as outlined in Remark 5.4 with $\beta=2$. We note that the set of Nash equilibria is unimpacted by these modifications. The Lipschitz constant of the cost functions over $\mathbb{X}$ is $L=7$. We apply Algorithm 4.8 with error levels $\varepsilon_{1}=\varepsilon_{2}=0.01$ to compute a set $X$ which satisfies (16) for $\epsilon=0.15$. This set $X$, along with the Nash equilibria and $\epsilon$-Nash equilibria, are displayed in Figure 5. We highlight that the computed set $X$ is comprised of distinct polyhedrons satisfying (16).
Figure 5: Example 5.5: the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ (yellow), the computed set $X$ (blue) and the set $\operatorname{NE}(f,\mathbb{X})$ (red).